# TRAIL: Trust-Aware Client Scheduling for Semi-Decentralized Federated Learning

### v1

@@ Introduction
This paper investigates semi-decentralized Federated Learning (SD-FL) as a framework to enhance the reliability of model training. As illustrated in Figure 1, we focus on a multi-edge server, multi-client SD-FL framework (Sun et al. 2021b). This architecture employs a two-tier aggregation approach: first, intra-cluster aggregation, where local models are aggregated by their respective servers, followed by inter-cluster consensus, where models from multiple servers are exchanged and collaboratively aggregated to train a shared global model. By distributing computational and communication loads, SD-FL improves both the robustness and scalability of the Federated Learning process. While SD-FL mitigates risks associated with single points of failure, existing research often overlooks the dynamic nature of clients, particularly fluctuations in model contributions and communication quality, which can adversely affect training efficiency (Sun et al. 2021a).
@@
Current research efforts primarily focus on the following areas. Communication Optimization: Efficient communication protocols are designed to reduce the frequency of interactions between clients and the server, lowering latency and bandwidth usage (Lin et al. 2021; Wang et al. 2024; Yemini et al. 2022; Sun et al. 2023). Trust Management: Trust mechanisms are introduced to evaluate and select clients for participation in training, thereby improving the overall credibility and performance of the model (Beltrán et al. 2023; Parasnis et al. 2023; Xu et al. 2024; Valdeira et al. 2023).
While there are studies that focus separately on communication optimization and trust management, existing methods often fail to effectively integrate these aspects. This gap negatively impacts the efficacy and efficiency of Federated Learning systems. An integrated approach that addresses both communication efficiency and participant reliability is essential for enhancing the robustness and performance of SD-FL, especially in diverse and potentially unreliable environments.
@@
$\begin{cases}\boldsymbol{b}_{i}\left(\boldsymbol{z}^{(s)}\right)=\boldsymbol{N}\left(\boldsymbol{Z};\boldsymbol{\mu}_{i}^{(s)}\boldsymbol{\Sigma}_{i}\right)\\ p_{i}(e)=N\left(e;\mu_{e}^{(s)},\sigma_{i}^{2}\right)\quad\\ \mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}\end{cases}.$ (9)
Here, $\boldsymbol{b}_{i}\left(\boldsymbol{z}^{(s)}\right)$ represents the probability density function for the state $i$ based on the observed data $\boldsymbol{z}^{(s)}$ from sensor $s$, modeled as a multivariate normal distribution with mean $\boldsymbol{\mu}_{i}^{(s)}$ and covariance $\boldsymbol{\Sigma}_{i}$. The term $p_{i}(e)$ signifies the probability density function for the dwell time $e$ in state $i$, also following a normal distribution characterized by mean $\mu_{e}^{(s)}$ and variance $\sigma_{i}^{2}$. The mean dwell time $\mu_{e}^{(s)}$ is calculated using the formula $\mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}$, where $\delta^{(s)}$ is a scaling factor, $m_{i}$ represents a parameter related to state $i$, and $\psi^{(s)}$ is a sensor-specific offset.
@@
$\begin{cases}B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi,\\ \Psi=\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i}<\Theta\right)\right)\right).\end{cases}$ (23)
@@
Input: Set of users $\mathcal{U}$, set of servers $\mathcal{S}$, maximum users per server $U_{\max}$
Output: Assignment matrix $\boldsymbol{d}$
@@ 2 Initialize client count $u_{s}=0,\forall s\in\mathcal{S}$;
3 Sort clients $i\in\mathcal{U}$ in increasing order based on $TL_{i}$ for each server $s$;
4 for each user $i\in\mathcal{U}$ do
5       for each server $s\in\mathcal{S}$ do
6             if $u_{s}<U_{\max}$ and $\sum_{\forall s\in\mathcal{S}}d_{i,s}=0$ then
7                   Assign user $i$ to server $s$: $d_{i,s}=1$;
8                   Increment user count for server $s$: $u_{s}=u_{s}+1$;
9                   break;
@@ 8                   Increment user count for server $s$: $u_{s}=u_{s}+1$;
    • Here, we have configured a system with five edge servers and fifty clients. Each client possesses 1,000 local training data samples. Additionally, we have set up the system so that approximately 30% of the clients experience a gradual decline in training quality as the training progresses.
@@
In Figure 2, we analyze the variations in test accuracy and training loss over multiple training rounds for four distinct mechanisms, specifically under conditions where only 10% of users are classified as low quality. Our proposed mechanism stands out by achieving the highest performance across four real datasets. This success can be attributed to the effective integration of our AHSMM and a greedy algorithm, which work synergistically to accurately predict fluctuations in client learning and communication quality. By optimizing the participation of low-quality clients, our approach significantly enhances overall training outcomes. In contrast, the TRUST mechanism, despite employing a predictive approach, lacks an effective client distribution strategy. This deficiency leads to suboptimal performance, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism incorporates the AHSMM to forecast client behavior, but it does not allocate clients efficiently, leading to less effective training sessions. Although the GUROBI mechanism is capable of determining an optimal client distribution scheme, its solution process is notably time-consuming, making it less efficient compared to our proposed greedy algorithm. Additionally, GUROBI does not incorporate client quality predictions, which hinders its ability to exclude low-quality clients from training participation in a timely manner, ultimately affecting the training efficiency. Overall, our mechanism demonstrates a superior ability to navigate the complexities of client quality by leveraging predictive modeling and strategic allocation, ensuring robust training performance in Federated Learning environments.
@@
In Figure 3, we analyze the comparative performance of four mechanisms across scenarios characterized by varying proportions of low-quality clients. As the percentage of low-quality clients increases, all mechanisms demonstrate a decline in both training and testing accuracy, albeit to different extents. Notably, our proposed mechanism consistently delivers the best results across four real datasets, showcasing its robustness in challenging environments. The superior performance of our mechanism can be attributed to the integration of the AHSMM and a greedy-based client allocation scheme. This combination effectively predicts fluctuations in client learning and communication quality, enabling efficient client allocation that significantly enhances model training quality, even under adverse conditions. By optimizing the participation of higher-quality clients, we mitigate the negative impact of low-quality users on overall performance. In contrast, while the TRUST mechanism is capable of identifying unreliable clients, it lacks an efficient client distribution strategy. This limitation results in poorer training outcomes compared to our mechanism, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism employs the AHSMM to accurately predict changes in client training quality but relies on a random allocation strategy during the client distribution phase. Consequently, this randomness undermines the final training effectiveness, leading to suboptimal results. The GUROBI mechanism, despite its potential for determining optimal client distributions, performs the worst in our experiments. Its inability to accurately predict changes in client learning quality restricts its effectiveness, as it cannot exclude low-quality clients in a timely manner. Our findings reveal a notable 8.7% increase in test accuracy and a 15.3% reduction in training loss compared to existing baselines, highlighting the efficacy of our approach in enhancing model performance in Federated Learning settings. These results underscore the importance of both predictive modeling and strategic client allocation in achieving high-quality training outcomes.

### v2

@@ Introduction
This paper investigates semi-decentralized Federated Learning (SD-FL) as a framework to enhance the reliability of model training. As illustrated in Figure 1, we focus on a multi-edge server and multi-client SD-FL framework (Sun et al. 2021b). This architecture employs a two-tier aggregation approach: first, intra-cluster aggregation, where local models are aggregated by their respective servers, followed by inter-cluster consensus, where models from multiple servers are exchanged and collaboratively aggregated to train a shared global model. By distributing computational and communication loads, SD-FL improves both the robustness and scalability of the Federated Learning process. While SD-FL mitigates risks associated with single points of failure, existing research often overlooks the dynamic nature of clients, particularly fluctuations in model contributions and communication quality, which can adversely affect training efficiency (Sun et al. 2021a).
@@
Current research efforts primarily focus on the following areas. Client Selection: Mechanisms are developed to select clients for participation in training, ensuring that chosen clients meet performance criteria, thereby enhancing the overall effectiveness of the model (Lin et al. 2021; Wang et al. 2024; Yemini et al. 2022; Sun et al. 2023). Trust Management: Trust mechanisms are introduced to assess and predict the reliability of clients, ensuring that only reliable clients participate in training. These mechanisms contribute to improved model robustness and performance (Beltrán et al. 2023; Parasnis et al. 2023; Xu et al. 2024; Valdeira et al. 2023).
While there are studies that focus separately on client selection and trust management, existing methods often fail to effectively integrate these aspects. This gap negatively impacts the efficacy and efficiency of SD-FL systems. An integrated approach that addresses both client selection and trust management is essential for enhancing the robustness and performance of SD-FL, especially in diverse and potentially unreliable environments.
@@
$\begin{cases}b_{i}\left(z^{(s)}\right)=N\left(Z;\mu_{i}^{(s)}\Sigma_{i}\right)\\ p_{i}(e)=N\left(e;\mu_{e}^{(s)},\sigma_{i}^{2}\right)\quad\\ \mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}\end{cases}.$ (9)
Here, $b_{i}\left(\boldsymbol{z}^{(s)}\right)$ represents the probability density function for the state $i$ based on the observed data $z^{(s)}$ from sensor $s$, modeled as a multivariate normal distribution with mean $\mu_{i}^{(s)}$ and covariance $\Sigma_{i}$. The term $p_{i}(e)$ signifies the probability density function for the dwell time $e$ in state $i$, also following a normal distribution characterized by mean $\mu_{e}^{(s)}$ and variance $\sigma_{i}^{2}$. The mean dwell time $\mu_{e}^{(s)}$ is calculated using the formula $\mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}$, where $\delta^{(s)}$ is a scaling factor, $m_{i}$ represents a parameter related to state $i$, and $\psi^{(s)}$ is a sensor-specific offset.
@@
$\begin{cases}B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi,\\ \Psi=\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i,s}<\Theta\right)\right)\right).\end{cases}$ (23)
@@
Input: Set of clients $\mathcal{U}$, set of servers $\mathcal{S}$, maximum clients per server $U_{\max}$, the allocation status $x_{i}$;
Output: Assignment matrix $\boldsymbol{d}$;
@@ 2 Initialize client count $u_{s}=0,\forall s\in\mathcal{S}$;
3 Sort $(i,s)\in(\mathcal{U},\mathcal{S})$ in increasing order based on $TL_{i,s}$ for each server $s$;
4 for (i,s) in $(\mathcal{U},\mathcal{S})$ do
5       if $u_{s}<U_{\max}$ and $x_{i}=0$ and $TL_{i,s}\geq\Theta$  then
6             Assign client $i$ to server $s$: $d_{i,s}=1$;
7             Increment client count for server $s$: $u_{s}=u_{s}+1$;
8             $x_{i}=1$;
9       end if
@@ 8             $x_{i}=1$;
    • We set up a system with five edge servers and fifty clients. Each client possesses 1,000 local training data samples. Additionally, we have set up the system so that 10%, 30%, and 50% of the clients experience a gradual decline in training quality and communication quality as the training progresses.
@@
In Figure 2, we analyze the variations in test accuracy and training loss over multiple training rounds for four distinct mechanisms, specifically under conditions where only 10% of clients are classified as low quality. Our proposed mechanism stands out by achieving the highest performance across four real datasets. This success can be attributed to the effective integration of our AHSMM and a greedy algorithm, which work synergistically to accurately predict fluctuations in client learning and communication quality. By optimizing the participation of low-quality clients, our approach significantly enhances overall training outcomes. In contrast, the TRUST mechanism, despite employing a predictive approach, lacks an effective client distribution strategy. This deficiency leads to suboptimal performance, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism incorporates the AHSMM to forecast client behavior, but it does not allocate clients efficiently, leading to less effective training sessions. Although the GUROBI mechanism can determine the optimal client allocation scheme, it does not incorporate client quality predictions, which hampers its ability to promptly exclude low-quality clients from participating in training, ultimately affecting training efficiency. Overall, our mechanism demonstrates a superior ability to navigate the complexities of client quality by leveraging predictive modeling and strategic allocation, ensuring robust training performance in Federated Learning environments.
@@
In Figure 3, we analyze the comparative performance of four mechanisms across scenarios characterized by varying proportions of low-quality clients. As the percentage of low-quality clients increases, all mechanisms demonstrate a decline in both training and testing accuracy, albeit to different extents. Notably, our proposed mechanism consistently delivers the best results across four real datasets, showcasing its robustness in challenging environments. The superior performance of our mechanism can be attributed to the integration of the AHSMM and a greedy-based client allocation scheme. This combination effectively predicts fluctuations in client learning and communication quality, enabling efficient client allocation that significantly enhances model training quality, even under adverse conditions. By optimizing the participation of higher-quality clients, we mitigate the negative impact of low-quality clients on overall performance. In contrast, while the TRUST mechanism is capable of identifying unreliable clients, it lacks an efficient client distribution strategy. This limitation results in poorer training outcomes compared to our mechanism, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism employs the AHSMM to accurately predict changes in client training quality but relies on a random allocation strategy during the client distribution phase. Consequently, this randomness undermines the final training effectiveness, leading to suboptimal results. The GUROBI mechanism, despite its potential for determining optimal client distributions, performs the worst in our experiments. Its inability to accurately predict changes in client learning quality restricts its effectiveness, as it cannot exclude low-quality clients in a timely manner. Our findings reveal a notable 8.7% increase in test accuracy and a 15.3% reduction in training loss compared to existing baselines, highlighting the efficacy of our approach in enhancing model performance in Federated Learning settings. These results underscore the importance of both predictive modeling and strategic client allocation in achieving high-quality training outcomes.